OpenStudy (anonymous):

Solve using the elimination method 5x+2y=-7 7x-3y=17

OpenStudy (asadkarim7):

3 ( 5x +2y =-7) 2 ( 7x -3y =17 ) 15x +6 y =- 21 14 x - 6 y = 34 eliminate 29 x = 13 x = 13 / 29

OpenStudy (anonymous):

this ones good

OpenStudy (anonymous):

multiply first by 7 and second by 5 to get 35x + 14y = -49 -----(3) 35x -15y = 85 -------- (4) subtract (4) from (3) to eliminate the variable x 29y = -134 y = -134/29 now put y= -134/29 in the first equation

OpenStudy (asadkarim7):

5 ( 13 /29 ) +2y = -7 2y = -7 - 65 / 39 2y = -273-65/39 y = 338 x 2 / 39 y = 676 /39

OpenStudy (anonymous):

5x + 2*(-134/29) = -7 5x -268/29 = -7 5x = -7 + 268/29 5x = 65/29 x = 65/145

OpenStudy (anonymous):

cancelling out common factors x = 13/29 so solution is x = 13/29 and y = -134/29

OpenStudy (anonymous):

@Asadkarim yr calculation for y is wrong and hence y= 676/39 is wrong.. correct value of y is -134/29

OpenStudy (asadkarim7):

yeah :(

OpenStudy (anonymous):

no problem, we all make mistakes every now n then.....

OpenStudy (anonymous):

rachelgreen, if u r satisfied with the answer, pls click GOOD ANSWER button for me :))

OpenStudy (anonymous):

dnt wrry buddy good one

OpenStudy (anonymous):

thanks vinit...........

OpenStudy (anonymous):

tats done